# Heuristic Methods are Good Teachers to Distill MLPs for Graph Link Prediction

### v1

@@ 2.3 GNN-to-MLP Distillation
3 Motivating Studies
In this section, we present key observations that motivate our use of heuristic methods for distilling MLPs. First, we compare GNN and GNN4LP models as teachers and find that GNN4LP, despite achieving better accuracy, does not always serve as a superior teacher. We then highlight the advantages of using heuristics as teachers: (1) they significantly enhance MLP performance and (2) they offer better training efficiency.
3.1 Better Models $\neq$ Better Teachers
Unlike node and graph classification tasks, state-of-the-art GNN4LP models for link prediction fall outside the scope of conventional message-passing neural networks (MPNNs). Since GNN4LP models usually outperform standard GNNs, one might intuitively expect GNN4LP-distilled MLPs to also surpass GNN-distilled MLPs. However, our findings contradict this assumption. As shown in Figure 1, while NCN [Wang et al., 2024] significantly outperforms SAGE [Hamilton et al., 2017] and GAT [Veličković et al., 2017], NCN-distilled MLPs exhibit suboptimal performance in three out of five datasets. This suggests that simply replacing GNNs with more powerful GNN4LP models does not necessarily enhance student’s performance.
Notably, prior studies have also observed that stronger teacher models do not always produce better students during knowledge distillation [Wu et al., 2024, Jafari et al., 2021, Zhu and Wang, 2021, Qiu et al., 2022]. They attribute it to the mismatch in capacity between teacher and student models. In our context, GNN4LP models leverage both node features and complex structural information, achieving strictly better capacity and expressiveness than standard GNN models [Zhang et al., 2021a]. In contrast, MLPs rely solely on node features and can only learn the structure information correlated with node features. Consequently, MLPs cannot fully inherit the effectiveness of GNN4LP models, leading to confusion and suboptimal performance due to guidance they are unable to interpret effectively.
The following theorem reveals why GNN4LP models are not bettet teachers than GNNs.
Theorem 3.1.
(GNN-distilled MLPs can be as good as any GNN4LP-distilled MLPs) Let $\mathcal{G}$ be a graph. Let $x_{i}$, $s_{i}$ denote the node features and latent structural information of node $i$. Assume $s_{i}=s_{i}^{\text{GNN}}+s_{i}^{\text{extra}}$, where $s_{i}^{\text{GNN}}$ is all the structural information capturable by GNNs from $\bm{x}_{i}$ and its $k$-hop neighbors, and $s_{i}^{\text{extra}}\perp\!\!\!\perp s_{i}^{GNN}$. $s_{i}^{extra}$ is the information that impossible for GNNs to capture, so we have $s_{i}^{\text{extra}}\perp\!\!\!\perp x_{i}$. A standard GNN can be written as $F_{\text{GNN}}(\bm{x}_{i},\bm{x}_{j},s_{i}^{\text{GNN}},s_{j}^{\text{GNN}})$. And since GNN4LP models enhance the model ability by incorporating information that impossible for GNN to capture, they can be written as $F_{\text{GNN4LP}}(\bm{x}_{i},\bm{x}_{j},s_{i}^{\text{GNN}},s_{j}^{\text{GNN}},s_{i}^{\text{extra}},s_{j}^{\text{extra}})$.
Notice that for any function $F$, the optimal MLP distilled from $F$ is $\mathbb{E}[F|x_{i},x_{j}]$. Then, for any $F_{\text{GNN4LP}}$, there exists a GNN $F_{\text{GNN}}$ such that the optimal MLP student distilled from both teachers satisfies:
$F_{\text{GNN4LP-MLP}}^{*}(\bm{x}_{i},\bm{x}_{j})=\mathbb{E}_{s^{\text{GNN}},s^{\text{extra}}}\left[F_{\text{GNN4LP}}\mid\bm{x}_{i},\bm{x}_{j}\right]$ (4)
$=$ $\mathbb{E}_{s^{\text{GNN}}}\left[F_{\text{GNN}}\mid\bm{x}_{i},\bm{x}_{j}\right]=F_{\text{GNN-MLP}}^{*}(\bm{x}_{i},\bm{x}_{j})$
The proof is given in Appendix C
(a) Hits@20 of teacher models
(b) Hits@20 of student MLPs
3.2 Heuristic Methods are good teachers
(a) Hits@10 of CN, MLP and CN-distilled MLP.
Since stronger GNN4LP models are not necessarily better teachers due to capacity mismatch, we then explore using simple heuristic methods as teachers. Unlike GNNs, heuristic methods typically capture a single type of structural information (e.g., common neighbors) [Mao et al., 2023], making them easier for MLPs to learn from. Additionally, heuristic methods can achieve accuracy comparable to, or even better than, GNN and GNN4LP models on certain datasets [Li et al., 2024], making them promising candidates for MLP distillation. However, their performance is highly data-dependent. On some datasets, heuristic methods may even perform worse than MLPs. Surprisingly, we find that even when heuristic methods underperform, they can still serve as effective teachers, and offering substantial training time savings compared to GNN-based teachers.
Observation 1: Heuristic methods are effective teachers for MLPs, even when they underperform.
Figure 2(a) presents the Hits@10 accuracy of CN, MLP, and CN-distilled MLP on three datasets. Although CN has lower accuracy than standard MLPs, the CN-distilled MLPs significantly outperform MLPs trained without distillation, indicating that CN-distilled MLPs effectively learn extra information during the distillation process. Additional results in Section 5 further confirm that heuristic-distilled MLPs achieve near-GNN performance across various datasets. These findings provide strong empirical evidence supporting the effectiveness of heuristic teachers.
Given the observation above, a natural question arises: why can heuristic methods improve MLPs even when they have lower accuracy? We hypothesize that heuristic methods provide complementary information to MLPs. Mao et al. [Mao et al., 2023] identify three fundamental factors critical to link prediction: local structure proximity, global structure proximity, and feature proximity. Different proximity capture distinct sets of positive node pairs in a graph. Specifically, they proposed a theoretical model for link prediction based on the underlying data factors:
$P(i\sim j)=\left\{\begin{aligned} \frac{1}{1+e^{\alpha(d_{ij}-\max\{r_{i},r_{j}\})}}&,d_{ij}\leq\max\{r_{i},r_{j}\}\\ \beta_{ij}&,d_{ij}>\max\{r_{i},r_{j}\}\end{aligned}\right.$ (5)
where $d_{ij}$ is the structural proximity between node $i,j$, and $\beta_{i,j}$ is the feature proximity. $r_{i}$ and $r_{j}$ are connecting threshold parameters for node $i$ and $j$. $\alpha>0$ controls the sharpness of the function. The theoretical model illustrates that when two nodes have close structure proximity, the likelihood of connection depends only on the structure proximity. Otherwise, the likelihood of connection only depends on the feature proximity.
Ma et al. [Ma et al., 2024] show that MLPs primarily capture feature proximity, exhibiting low correlation with structure proximity, an intuitive result since MLPs rely solely on node features. So MLPs primarily focus on node pairs where $d_{ij}>\max\{r_{i},r_{j}\}$ in the theoretical model (Eq. 5). In contrast, heuristics such as CN, AA, and RA capture local structure proximity, while CSP captures global structure proximity [Mao et al., 2023]. As a result, heuristic methods identify positive node pairs where $d_{ij}\leq\max\{r_{i},r_{j}\}$, the pairs MLPs alone would overlook. In addition, since each heuristic only captures one type of structure proximity, it is easier for MLP to learn the correlation between node features and the structural proximity, enabling them to capture node pairs beyond the scope of feature proximity.
To validate our hypothesis, we compare the positive edges identified by heuristic methods, MLPs, and heuristic-distilled MLPs. Following the approach in [Ma et al., 2024], we use Hits@K (K=5 for Citeseer, K=10 for Cora and Pubmed) to identify the sets of positive edges captured by different methods. Specifically, given a positive edge and a set of negative edges, if the positive edge ranks among the top K highest predicted values, it is included in the positive edge set. Once the positive edge sets are identified, we compute the subset ratio, which represents the proportion of positive edges identified by heuristic methods that are also recognized by MLPs. Figure 3 presents the results. As shown, distilling from heuristic methods significantly increases the subset ratio, indicating that heuristic-distilled MLPs effectively capture extra positive edges and learn complementary information from heuristic teachers.
Next, we offer the following theoretical perspective to explain when and why heuristics can be at least as effective, or even better than GNNs as teachers.
Previous studies [Ma et al., 2024, Mao et al., 2023] have shown that the primary challenge of training MLPs lies in learning the structural term, i.e., the node pairs where $d_{ij}\leq\max\{r_{i},r_{j}\}$. Let $\tilde{f}(x_{i},x_{j},s_{i},s_{j})$ be a GNN or heuristic that approximates $f(i,j)=\frac{1}{1+e^{\alpha(d_{ij}-\max\{r_{i},r_{j}\})}}$. The best MLP distilled from $\tilde{f}$ is $\mathbb{E}\tilde{f}=\mathbb{E}\left(\tilde{f}(x_{i},x_{j},s_{i},s_{j})|x_{i},x_{j}\right)$. Consider the KL Divergence
$KL(f\|\mathbb{E}\tilde{f})$
$=$ $-\mathbb{E}_{x_{i},x_{j},s_{i},s_{j}}\left(\sum f\log\frac{f}{\mathbb{E}\tilde{f}}\right)$
$=$ $-\mathbb{E}_{x_{i},x_{j},s_{i},s_{j}}\left(\sum f\log\frac{f*\tilde{f}}{\tilde{f}*\mathbb{E}\tilde{f}}\right)$
$=$ $KL(f\|\tilde{f})+\mathbb{E}_{x_{i},x_{j},s_{i},s_{j}}\mathbb{E}_{f}(\log\mathbb{E}\tilde{f}-\log\tilde{f})$
This decomposition shows that the total distillation error consists of:
  1. 1.
    The teacher’s error: $KL(f\|\tilde{f})$, i.e., how well the teacher approximates the true label distribution $f$;
  2. 2.
    The student’s approximation error: how well the student (MLP) can recover $\tilde{f}$ from features alone.
Although GNNs may have lower $KL(f\|\tilde{f})$, their greater complexity can lead to larger student error. In contrast, heuristic methods may produce MLP students with lower total KL divergence due to better alignment with the student’s capacity.
Theorem 3.2 (Sufficient Condition for Heuristics to Be Better Teachers).
Let $h_{ij}$ be a heuristic teacher and $\tilde{f}_{ij}$ a GNN teacher. If
$KL(f\|h)+\mathbb{E}_{f}\left[\log\mathbb{E}(h_{ij}\mid x_{i},x_{j})-\log h_{ij}\right]$
$\leq$ $KL(f\|\tilde{f})+\mathbb{E}_{f}\left[\log\mathbb{E}(\tilde{f}_{ij}\mid x_{i},x_{j})-\log\tilde{f}_{ij}\right],$
then the optimal MLP distilled from $h$ achieves lower KL divergence with respect to $f$ than the MLP distilled from $\tilde{f}$.
This theorem highlights that a simpler heuristic teacher can outperform a more expressive GNN teacher in the context of distillation, when the student cannot effectively learn the complex function class of the teacher.
Observation 2: Heuristic methods are much faster than GNNs.
Despite their effectiveness, GNNs and GNN4LP models suffer from high computational costs during both training and inference. Figure 2(b) illustrates the time breakdown of GNN-to-MLP distillation on three large-scale datasets: OGBL-Collab, OGBL-Citation2 [Hu et al., 2020], and IGB [Khatua et al., 2023]. The total GNN time includes both GNN training and guidance generation, while the MLP time refers to the MLP distillation process based on GNN guidance. We observe that 62.2% to 87.1% of the total time is spent on GNNs, making it the primary bottleneck in the distillation pipeline.
4 Ensemble Heuristic-Distilled MLPs
Section 3.2 highlights the advantages of using heuristic methods as teachers. However, since each heuristic method captures only a specific type of structural proximity [Mao et al., 2023], an MLP distilled from a single heuristic is inherently limited, capturing only a subset of positive links. To improve effectiveness, we aim to develop an approach that enables MLPs to leverage multiple heuristics.
4.1 Pitfall of Ensemble Heuristics
Ma et al. [Ma et al., 2024] demonstrated that combining multiple heuristics generally improves performance. So a straightforward approach is to first ensemble multiple heuristics and then distill a single MLP from the combined heuristic signals. However, as discussed in Section 3.1, a stronger model does not always translate into a better teacher. Unfortunately, this holds true when mixing heuristics.
Figure 4 presents the accuracy rankings (Hits@20) of CN, CSP, and their combination (CN+CSP), alongside the accuracy rankings of their corresponding student MLPs. While the combined heuristic indeed achieves higher accuracy, its distilled MLP performs worse than those distilled from individual heuristics. We hypothesize that this occurs because mixing heuristics increases structural complexity, making it more difficult for MLPs to learn effectively. These results indicate that directly distilling from a mixture of heuristics is not a viable solution.
4.2 Ensemble Heuristic-Distilled MLPs
An alternative approach is to ensemble multiple MLPs. For example, AdaGMLP [Lu et al., 2024] proposes a boosting framework to distill multiple MLPs from a GNN for node classification. However, AdaGMLP is not directly applicable to link prediction or scenarios involving multiple heuristic teachers. Moreover, boosting requires iterative training of multiple MLPs, which can be slow.
To effectively and efficiently ensemble multiple MLPs, our key insight is that: Since different heuristics capture different sets of positive links, their distilled MLPs should also capture different sets of positive links. To validate this, we compute the overlapping ratio of positive edges identified by different distilled MLPs. We use Hits@10 to define the positive edge set for each distilled MLP. More details about positive edge set can be found in Section 3.2. The overlapping ratio is then computed as the Jaccard index between two positive edge sets. As shown in Figure 5, the overlap among different heuristic-distilled MLPs is relatively low, confirming our hypothesis that each MLP captures complementary link information.
So a simple yet effective solution is to train a gating function that dynamically selects which heuristic-distilled MLP to use for each input node pair. Ma et al. [Ma et al., 2024] introduced a similar idea by training a gating network to ensemble different GNN4LP models. However, their gating network is not suitable for our setting, as it takes heuristic values as inputs, requiring neighbor-fetching operations, which would compromise the inference efficiency of MLPs.
To train the gating function, we fix the parameters of the distilled MLPs and optimize the parameters of the gating function using the BCE loss:
$-y_{ij}\log(\sum_{h\in\mathcal{H}}w_{h}q^{(h)}_{i,j})-(1-y_{ij})\log(1-\sum_{h\in\mathcal{H}}w_{h}q^{(h)}_{i,j})$ (6)
where $y_{ij}$ is the ground-truth label, $\mathcal{H}$ is the heuristic set, and $q^{(h)}$ is the prediction of the MLP distilled from heuristic $h$, and $w_{h}$ is the output weight of the gating function.
Additionally, we incorporate L1-regularization on the gate weights to encourage sparsity and improve generalization. The final loss function for training the gating MLP is:
$L_{BCE}+\lambda\sum_{h\in\mathcal{H}}|w_{h}|$ (7)
where $\lambda$ is a hyper-parameter controlling the weight of L1-regularization.
(a) Cora
(b) Pubmed
(c) Coauthor-CS
Figure 5: Overlap ratio of student MLPs distilled from different heuristic methods across datasets.
Figure 6: Illustration of Ensemble Heuristic-Distilled MLPs.
5 Experiments
In this section we introduce our main experiment results. Additional experiment details and ablation studies can be found in the appendix.
Dataset MLP Best GNN CN AA RA CSP
Cora 73.21$\pm$3.28 73.96$\pm$1.23 42.69$\pm$0.00 42.69$\pm$0.00 42.69$\pm$0.00 42.69$\pm$0.00
Citeseer 68.26$\pm$1.92 85.10$\pm$2.25 35.16$\pm$0.00 35.16$\pm$0.00 35.16$\pm$0.00 58.02$\pm$0.00
Pubmed 49.40$\pm$3.53 68.83$\pm$2.84 27.93$\pm$0.00 27.93$\pm$0.00 27.93$\pm$0.00 27.93$\pm$0.00
CS 39.36$\pm$0.99 64.24$\pm$3.52 53.65$\pm$0.00 74.32$\pm$0.00 74.01$\pm$0.00 0.00$\pm$0.00
Physics 21.60$\pm$3.09 65.85$\pm$2.82 61.40$\pm$0.00 78.10$\pm$0.00 79.80$\pm$0.00 0.00$\pm$0.00
Computers 17.53$\pm$1.25 27.07$\pm$4.45 20.38$\pm$0.00 27.54$\pm$0.00 32.67$\pm$0.00 0.00$\pm$0.00
Photos 31.81$\pm$2.82 49.79$\pm$8.87 34.37$\pm$0.00 42.63$\pm$0.00 45.39$\pm$0.00 0.00$\pm$0.00
Collab 44.38$\pm$3.47 59.14$\pm$1.64 61.37$\pm$0.00 64.17$\pm$0.00 63.81$\pm$0.00 46.49$\pm$0.00
IGB 19.13$\pm$1.34 20.47$\pm$1.39 7.78$\pm$0.00 7.78$\pm$0.00 7.78$\pm$0.00 1.00$\pm$0.00
Citation2 39.17$\pm$0.44 84.90$\pm$0.06 74.30$\pm$0.00 75.96$\pm$0.00 76.04$\pm$0.00 0.28$\pm$0.00
Dataset MLP Best LLP CN MLP AA MLP RA MLP CSP MLP
Cora 73.21$\pm$3.28 76.62$\pm$2.00 75.75$\pm$2.45 74.84$\pm$3.48 75.94$\pm$4.35 74.16$\pm$4.52
Citeseer 68.26$\pm$1.92 76.53$\pm$4.88 75.69$\pm$1.14 70.90$\pm$2.24 75.38$\pm$1.36 77.10$\pm$2.10
Pubmed 49.40$\pm$3.53 59.95$\pm$1.58 56.98$\pm$2.47 54.95$\pm$2.91 60.21$\pm$3.71 57.82$\pm$3.36
CS 39.36$\pm$0.99 66.64$\pm$1.80 67.53$\pm$5.26 69.56$\pm$2.97 69.39$\pm$2.97 67.41$\pm$1.34
Physics 21.60$\pm$3.09 60.19$\pm$2.93 57.09$\pm$4.13 61.10$\pm$3.15 52.87$\pm$4.12 50.80$\pm$2.09
Computers 17.53$\pm$1.25 25.80$\pm$4.80 27.55$\pm$2.88 30.45$\pm$4.27 25.50$\pm$1.97 21.97$\pm$1.26
Photos 31.81$\pm$2.82 39.66$\pm$2.94 39.17$\pm$4.18 40.23$\pm$4.07 30.84$\pm$3.25 38.49$\pm$5.06
Collab 44.38$\pm$3.47 49.30$\pm$0.79 48.23$\pm$0.89 47.33$\pm$1.02 48.93$\pm$0.66 48.99$\pm$0.69
IGB 19.13$\pm$1.34 25.12$\pm$1.14 24.38$\pm$0.11 24.20$\pm$0.90 24.11$\pm$1.25 24.75$\pm$0.83
Citation2 39.17$\pm$0.44 42.78$\pm$0.10 43.05$\pm$0.23 43.30$\pm$0.08 42.90$\pm$0.12 43.17$\pm$0.11
5.1 Experiment Setting
Datasets. Following previous works [Guo et al., 2023], we include ten datasets for a comprehensive evaluation, i.e., Cora, Citeseer, Pubmed [Yang et al., 2016], Coauthor-CS, Coauthor-Physics [Shchur et al., 2018], Amazon-Computers, Amazon-photos [McAuley et al., 2015, Shchur et al., 2018], OGBL-Collab, OGBL-Citation2 [Wang et al., 2020], and IGB [Khatua et al., 2023]. Dataset statistics are summarized in Table 4 in the appendix. Notably, OGBL-Collab contains over a million edges, while IGB and OGBL-Citation2 have over a million nodes and tens of millions of edges.
Evaluation Protocol. For Collab and Citation2 datasets, we use the same train/test split as in the OGBL benchmark [Hu et al., 2020], where the edges are split according to time to simulate real-world production settings. For all other datasets, we randomly sample 5%/15% of the edges along with an equal number of non-edge node pairs, as validation/test sets. Unlike LLP [Guo et al., 2023], which trains the teacher GNN with ten different random seeds and selects the model with the highest validation accuracy, we train the teacher GNN only once for distillation. Our approach better reflects practical constraints, as training a GNN ten times is infeasible for real-world applications. For each dataset, we train three teacher GNNs: GCN [Kipf and Welling, 2016a], SAGE [Hamilton et al., 2017], and GAT [Veličković et al., 2017], and report the highest-performing one, denoted as “Best GNN”. Similarly, for the LLP baseline [Guo et al., 2023], we distill three MLPs from GCN, SAGE, and GAT, respectively, and report the best-performing one, denoted as “Best LLP”. We include four heuristic methods: CN [Newman, 2001], AA [Adamic and Adar, 2003], RA [Zhou et al., 2009], and CSP [Liben-Nowell and Kleinberg, 2003]. For each dataset, all MLPs have same architecture. For Collab we report Hits@50, for IGB we report Hits@100, for Citation2 we report MRR, for other datasets we report Hits@20. More details of the hyper-parameters and breakdowns of “Best GNN” and “Best LLP” are reported in the appendix.
Dataset MLP Best GNN Best LLP EHDM $\Delta_{\text{GNN}}$ $\Delta_{\text{LLP}}$
cora 73.21$\pm$3.28 73.96$\pm$1.23 76.62$\pm$2.00 80.49$\pm$1.51 +8.83% +5.06%
citeseer 68.26$\pm$1.92 85.10$\pm$2.25 76.53$\pm$4.88 79.08$\pm$1.90 -7.08% +3.33%
pubmed 49.40$\pm$3.53 68.83$\pm$2.84 59.95$\pm$1.58 60.75$\pm$3.42 -11.74% +1.33%
cs 39.36$\pm$0.99 64.24$\pm$3.52 66.64$\pm$1.80 74.33$\pm$2.59 +15.71% +11.53%
physics 21.60$\pm$3.09 65.85$\pm$2.82 60.19$\pm$2.93 64.44$\pm$5.22 -2.14% +7.06%
computers 17.53$\pm$1.25 27.07$\pm$4.45 25.80$\pm$4.80 30.41$\pm$2.90 +12.35% +17.89%
photos 31.81$\pm$2.82 49.79$\pm$8.87 39.66$\pm$2.94 45.89$\pm$1.67 -7.83% +15.71%
collab 44.38$\pm$3.47 59.14$\pm$1.64 49.30$\pm$0.79 49.27$\pm$0.88 -16.70% -0.07%
IGB 17.74$\pm$1.20 20.47$\pm$0.82 25.12$\pm$1.14 27.27$\pm$0.27 +33.21% +8.54%
citation2 38.12$\pm$0.18 84.90$\pm$0.06 42.78$\pm$0.10 46.58$\pm$0.10 -45.13% +8.87%
5.2 How do Heuristic-Distillation Compare to GNN-Distillation
First, we provide more empirical results to show the effectiveness of heuristic methods as teachers for MLP distillation. Table 1 compares the teacher performance of GNNs and heuristic methods, while Table 2 reports the performance of their corresponding student models across all datasets. We observe that despite heuristic methods having significantly lower performance on certain datasets, heuristic-distilled MLPs consistently match or surpass the performance of the best GNN-distilled MLPs. These results empirically demonstrate that heuristic methods can serve as effective teachers for training MLPs.
5.3 Performance of Ensemble Approach
Figure 7: Training time decomposition of SAGE-LLP and our ensemble method.
Table 3 compares the performance of MLP, GNN, LLP (GNN-distilled MLP), and our EHDM, which ensembles four heuristic-distilled MLPs. The results show that EHDM outperforms both MLP and LLP by leveraging the strengths of different heuristic-distilled MLPs. It is on average $7.93\%$ better than the best LLP model. Notably, it achieves at least 90% of the best GNN effectiveness on 7 out of 10 datasets and even surpasses the best GNN on 4 datasets. For the two datasets where EHDM and LLP perform suboptimally (i.e., Collab and Citation2), we attribute this to their simple node features, as both datasets contain only 128-dimensional node embeddings, the smallest among all datasets.
Additionally, Figure 7 illustrates the training time decomposition of our ensemble method compared to the best LLP. The ensemble pipeline consists of three stages: heuristic guidance computation, MLP distillation, and ensembling. Although heuristic guidance must be generated for multiple heuristics, and MLP distillation is performed separately for each heuristic, both of these stages run in parallel. Therefore, we take the maximum time among the four heuristics as the effective time cost for each of these stages. From the figure, we observe that the ensemble time, corresponding to training the gating MLPs, constitutes only a small fraction of the overall pipeline. More importantly, by eliminating GNN training, our ensemble approach accelerates the entire distillation process by 1.95–3.32$\times$ compared to LLP.
6 Conclusion
In this paper,we demonstrate that simple heuristic methods, despite their lower accuracy, can be surprisingly effective teachers for MLPs, enabling competitive link prediction performance while drastically reducing training costs. Furthermore, we introduce an ensemble approach that aggregates multiple heuristic-distilled MLPs using a gating mechanism. Our extensive experiments across ten datasets show that this approach consistently improves prediction accuracy, while saving overall training time. Our findings have important implications for real-world link prediction tasks, particularly in large-scale web applications
Supplementary Material
Appendix A Hyper-parameters

### v2

@@ Appendix A LLP framework
Appendix B Proof of Lemma 3.1
Proof.
Expanding the first objective using the definition of KL divergence:
$\mathbb{E}_{x_{i},x_{j},s_{i},s_{j}}\left[\mathrm{KL}(F\,\|\,g)\right]$
$=$ $\mathbb{E}_{x_{i},x_{j},s_{i},s_{j}}\left[\sum_{y}F(y\mid x_{i},x_{j},s_{i},s_{j})\log\frac{F(y\mid x_{i},x_{j},s_{i},s_{j})}{g(y\mid x_{i},x_{j})}\right]$
$=$ $\mathbb{E}_{x_{i},x_{j}}\left[\sum_{y}\left(\mathbb{E}_{s_{i},s_{j}\mid x_{i},x_{j}}[F(y\mid x_{i},x_{j},s_{i},s_{j})]\cdot\log\frac{\mathbb{E}_{s_{i},s_{j}\mid x_{i},x_{j}}[F(y\mid\cdot)]}{g(y\mid x_{i},x_{j})}\right)\right]+C$
$=$ $\mathbb{E}_{x_{i},x_{j}}\left[\mathrm{KL}\left(\mathbb{E}[{F}|x_{i},x_{j}]\,\|\,g(y\mid x_{i},x_{j})\right)\right]+C,$
where $C$ is a constant independent of $g$, arising from the entropy terms of $F$.
Meanwhile, since $\mathbb{E}[F|x_{i},x_{j}]$ are $g$ are both independent to $s_{i},s_{j}$, so
$\mathbb{E}_{x_{i},x_{j},s_{i},s_{j}}KL(\mathbb{E}[F|x_{i},x_{j}],g)=\mathbb{E}_{x_{i},x_{j}}KL(\mathbb{E}[F|x_{i},x_{j}],g)$
Therefore, the two objectives are equivalent. ∎
Appendix C Proof of Theorem 3.2
Proof.
The teachable knowledge of $F_{\text{GNN4LP-MLP}}$ is
$\mathbb{E}_{s^{\text{GNN}},s^{\text{extra}}}\left[F_{\text{GNN4LP}}\mid\bm{x}_{i},\bm{x}_{j}\right]$ (11)
Let a GNN function be
$F_{\text{GNN}}(\bm{x}_{i},\bm{x}_{j},s_{i}^{\text{GNN}},s_{j}^{\text{GNN}})=\mathbb{E}[F_{GNN4LP}|\bm{x}_{i},\bm{x}_{j},s_{i},s_{j}]$
Then its teachable knowledge
$\mathbb{E}[F_{GNN}|\bm{x}_{i},\bm{x}_{j}]=\mathbb{E}\{\mathbb{E}[F_{GNN4LP}|\bm{x}_{i},\bm{x}_{j},s_{i},s_{j}]\mid\bm{x}_{i},\bm{x}_{j}\}$
With the tower property of the expectation, i.e. $\mathbb{E}(f|X)=\mathbb{E}(\mathbb{E}(f|X,Y)|X)$, we know that
$\mathbb{E}[F_{GNN}|\bm{x}_{i},\bm{x}_{j}]=\mathbb{E}[F_{GNN4LP}|\bm{x}_{i},\bm{x}_{j}]$ (12)
Therefore we prove that for any GNN4LP we can find a GNN teacher to have the same teachable knowledge. ∎
Appendix D Computation Steps of Toy Example
Let $x_{i},x_{j},s_{i},s_{j}$ all be binary random variables. And $Y_{ij}=1$ if and only if $s_{i}=s_{j}=1$. Let $x_{i},x_{j}\overset{iid}{\sim}\text{Bern}(0.8)$. Let $s|x\sim\text{Bern}(p_{x})$, with
$p_{x}=\left\{\begin{aligned} &0.5,\quad x_{i}=x_{j}=1,\\ &0.6.\quad\text{otherwise},\end{aligned}\right$. (13)
Let us consider two teacher functions. The first teacher $F_{1}(x_{i},x_{j},s_{i},s_{j})=1$ if and only if $s_{i}=s_{j}=1$, and $F_{1}=0.1$ otherwise. The second teacher $F_{2}$ is a constant predictor and $F_{2}=0.3$. So the teachable knowledge of $F_{1}$ is $\mathbb{E}[F_{1}|x_{i},x_{j}]=0.1+0.9p_{x}^{2}$. And the teachable knowledge of $F_{2}$ is $\mathbb{E}[F_{2}|x_{,}x_{j}]=0.3$.
[…]
$\mathbb{E}\bigl[-\ln q(Y)\mid x\bigr]=p_{x}^{2}\bigl(-\ln q\bigr)\;+\;(1-p_{x}^{2})\bigl(-\ln(1-q)\bigr).$
The following computation steps are illustrated in Table 4 and Table 5.
Table 4: Conditional expected NLL given $x$. Here $p_{x}=0.5$ if $(x_{i},x_{j})=(1,1)$ (prob. 0.64), else $p_{x}=0.6$ (prob. 0.36).
Region Predictor $q$ $\mathbb{E}[-\ln q(Y)\mid x]$
$(1,1)$, $p_{x}^{2}=0.25$ $F_{1}$ $q_{1}=1\,(Y=1),\,0.1\,(Y=0)$ $(1-0.25)(-\ln 0.9)\approx 0.0790$
$F_{2}$ $q_{2}=0.3$ $0.25(-\ln 0.3)+0.75(-\ln 0.7)\approx 0.5685$
$\mathbb{E}F_{1}$ $q_{g1}=0.1+0.9\cdot 0.25=0.325$ $\approx 0.5758$
$\mathbb{E}F_{2}$ $q_{g2}=0.3$ $\approx 0.5685$
otherwise, $p_{x}^{2}=0.36$ $F_{1}$ $q_{1}=0.1$ $(1-0.36)(-\ln 0.9)\approx 0.0674$
$F_{2}$ $q_{2}=0.3$ $0.36(-\ln 0.3)+0.64(-\ln 0.7)\approx 0.6617$
$\mathbb{E}F_{1}$ $q_{g1}=0.1+0.9\cdot 0.36=0.424$ $\approx 0.6619$
$\mathbb{E}F_{2}$ $q_{g2}=0.3$ $\approx 0.6617$
Table 5: Unconditional expected NLL: $\Pr((1,1))=0.64,\ \Pr(\text{otherwise})=0.36$.
Predictor $\mathbb{E}[-\ln q(Y)]$
$F_{1}$ $0.64\cdot 0.0790+0.36\cdot 0.0674\approx 0.0748$
$F_{2}$ $0.64\cdot 0.5685+0.36\cdot 0.6617\approx 0.6021$
$\mathbb{E}F_{1}$ $0.64\cdot 0.5758+0.36\cdot 0.6619\approx 0.6068$
$\mathbb{E}F_{2}$ same as Teacher 2: $\approx 0.6021$
Appendix E Experiment Setting
Datasets. Following previous works [guo2023linkless], we include ten datasets for a comprehensive evaluation, i.e., Cora, Citeseer, Pubmed [yang2016revisiting], Coauthor-CS, Coauthor-Physics [shchur2018pitfalls], Amazon-Computers, Amazon-photos [mcauley2015image, shchur2018pitfalls], OGBL-Collab, OGBL-Citation2 [wang2020microsoft], and IGB [khatua2023igb]. Dataset statistics are summarized in Table 6 in the appendix. Notably, OGBL-Collab contains over a million edges, while IGB and OGBL-Citation2 have over a million nodes and tens of millions of edges.
Evaluation Protocol. For Collab and Citation2 datasets, we use the same train/test split as in the OGBL benchmark [ogb], where the edges are split according to time to simulate real-world production settings. For all other datasets, we randomly sample 5%/15% of the edges along with an equal number of non-edge node pairs, as validation/test sets. Unlike LLP [guo2023linkless], which trains the teacher GNN with ten different random seeds and selects the model with the highest validation accuracy, we train the teacher GNN only once for distillation. Our approach better reflects practical constraints, as training a GNN ten times is infeasible for real-world applications. For each dataset, we train three teacher GNNs: GCN [kipf2016semi], SAGE [hamilton2017inductive], and GAT [velivckovic2017graph], and report the highest-performing one, denoted as “Best GNN”. Similarly, for the LLP baseline [guo2023linkless], we distill three MLPs from GCN, SAGE, and GAT, respectively, and report the best-performing one, denoted as “Best LLP”. We include four heuristic methods: CN [newman2001clustering], AA [adamic2003friends], RA [zhou2009predicting], and CSP [liben2003link]. For each dataset, all MLPs have same architecture. For Collab we report Hits@50, for IGB we report Hits@100, for Citation2 we report MRR, for other datasets we report Hits@20.
Appendix F Hyper-parameters